# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v2

@@ 1.4. Proof Strategy and Paper Organization
We prove Theorem 1 by the strategy illustrated in Fig. 1.1. Using tools from optimal transport [47], we first show in Section 2 that the nonlinear log-Sobolev inequality implies the nonlinear transport inequality. By our previous result [48] from information theory, we then show in Section 3 that the nonlinear transport inequality implies a variant isoperimetric inequality—the isoperimetric inequality with thick boundaries. Lastly, by tools from geometric measure theory [34], we show in Section 4 that the variant isoperimetric inequality implies the original isoperimetric inequality. Combining these three steps yields that the nonlinear log-Sobolev inequality implies the isoperimetric inequality. As for the other direction, we provide two proofs in Section 5: one is based on information theory and optimal transport, and the other is based on geometric measure theory.
We prove Theorem 2, Theorem 4, and Proposition 3 respectively in Sections 6-9.
@@ 2. From Log-Sobolev to Transport
2.1. Connection between Log-Sobolev and Transport
@@
Let $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ be a probability measure with density $h=\frac{{\text{\rm d}}\mu}{{\text{\rm d}}\nu}$ such that
@@
Define a Laplace-type operator $L=\Delta-\nabla V\cdot\nabla$. The heat equation $\partial_{t}\rho=L\rho$ with initial data $\rho_{0}=h$ admits a solution $\rho_{t}$; see [38]. Let $\mu_{t}=\rho_{t}\nu$. The heat equation has intimate relations with the entropy and the Wasserstein metric in the sense that the heat equation is the gradient flow of the entropy functional $D$ with respect to the Wasserstein metric $\mathsf{W}_{2}$.
It is already known that Sobolev-type inequalities (including log-Sobolev inequalities) imply Talagrand’s transport inequalities; this point was shown by Otto and Villani [38]. We now show a nonlinear analogue of this result. That is, nonlinear log-Sobolev inequalities imply a nonlinear version of Talagrand’s transport inequalities.
Theorem 5 (From Nonlinear Log-Sobolev to Nonlinear Transport).
Assume that holds the inequality
for some function $\theta$. Then, for $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ satisfying (2.1), the solution $\mu_{t}$ given above satisfies
(2.2) $\mathsf{W}_{2}\left(\mu,\mu_{t}\right)\leq\int_{D(\mu_{t}\|\nu)}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
In particular, letting $t\to\infty$ yields that
(2.3) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\int_{0}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Proof.
The proof follows Otto and Villani’s proof idea of Theorem 1 in [38]. By Lemmas 1 and 2 in [38],
$\frac{{\text{\rm d}}}{{\text{\rm d}}t}D(\mu_{t}\|\nu)$ $=-I(\mu_{t}\|\nu),$
$\frac{{\text{\rm d}}^{+}}{{\text{\rm d}}t}\mathsf{W}_{2}\left(\mu,\mu_{t}\right)$ $\leq\sqrt{I(\mu_{t}\|\nu)}\textrm{ for almost all }t,$
where $\frac{{\text{\rm d}}^{+}}{{\text{\rm d}}t}$ stands for the upper right derivative.
Define
$f(t)=\mathsf{W}_{2}\left(\mu,\mu_{t}\right)+\Upsilon(D(\mu_{t}\|\nu)),$
where
$\Upsilon(s)=\int_{0}^{s}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Then, it holds that
$\frac{{\text{\rm d}}^{+}}{{\text{\rm d}}t}f(t)$ $\leq\sqrt{I(\mu_{t}\|\nu)}+\frac{\frac{{\text{\rm d}}}{{\text{\rm d}}t}D(\mu_{t}\|\nu)}{\sqrt{\theta(D(\mu_{t}\|\nu))}}$
$=\sqrt{I(\mu_{t}\|\nu)}-\frac{I(\mu_{t}\|\nu)}{\sqrt{\theta(D(\mu_{t}\|\nu))}}$
$\leq 0.$
So, $f$ is non-increasing as a function of $t$, which implies $f(t)\leq f(0)=\Upsilon(D(\mu\|\nu))$ for any $t\geq 0$. This proves Theorem 5. ∎
Theorem 6 (From Nonlinear Log-Sobolev to Nonlinear Transport).
Assume that holds the inequality
$\theta(D(\lambda\|\nu))\leq I(\lambda\|\nu),\;\forall\lambda\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
for some function $\theta$. Then, for any $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ (which does not necessarily satisfy (2.1)) and any $0\leq s\leq D(\mu\|\nu)$, there is a probability measure $\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\pi\|\nu)=s$ and
In particular, letting $s=0$ yields that
(2.5) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\int_{0}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
If $\theta$ is set to the linear function $\theta(r)=2Kr$, then the nonlinear transport inequality in (2.5) reduces to the classic Talagrand transport inequality.
Proof.
If $\mu$ satisfies (2.1), this theorem follows directly from Theorem 5 since $D(\mu_{t}\|\nu)$ is continuous in $t$. We next consider the case that $\mu$ does not satisfy (2.1).
It was shown in [38] that for every $\mu\in P_{2}^{\mathrm{ac}}(M)$, there is a sequence of probability measures $\mu_{k}$ satisfying (2.1) such that as $k\to\infty$,
@@
Given any $0<s<D(\mu\|\nu)$, it holds that $0<s<D(\mu_{k}\|\nu)$ for all sufficiently large $k$. By Theorem 5, for each such $k$, there is a probability measure $\pi_{k}$ such that $D(\pi_{k}\|\nu)=s$ and
@@
Note that (2.7) implies that $\mu_{k}$ belong to a sublevel set of the relative entropy $\pi\mapsto D(\pi\|\nu)$, and the sublevel set of the relative entropy is weakly compact. So, $\pi_{k}$ converges to some $\pi^{*}$ weakly. Since $\mu$ has finite second moment, (2.6) implies that the sequence $\mu_{k}$ has bounded second moment. In this case, $\pi_{k}\to\pi^{*}$ weakly is equivalent to $\mathsf{W}_{2}\left(\pi_{k},\pi^{*}\right)\to 0$. Taking limits for (2.8) and applying the triangle inequality yield
@@
The importance of Theorems 5 and 6 is as follows. Given $\tau\geq 0,\mu,\nu$, consider the following optimization problem:
@@ 2.2. Generalization to $\Phi$-Sobolev
Theorems 5 and 6 can be generalized to the general $\Phi$-entropy setting (also known as $U$-functional or $f$-divergence). Let $\Phi:[0,+\infty)\rightarrow\mathbb{R}$ be a continuous convex function, twice differentiable on $(0,+\infty)$. The $\Phi$-entropy is defined as
@@
A common requirement for the function $\Phi$ is the following, which guarantees the displacement convexity of the $\Phi$-entropy [47, Chapter 16].
@@
Based on this theorem, an analogue of Theorem 6 for $\Phi$-entropy can be established as well.
@@ 3. From Transport to Variant Isoperimetry
In this section, we show how entropy-transport inequalities can be applied to obtain isoperimetric inequalities. We divide this part into two settings: the isoperimetric problem with thick boundary and the standard isoperimetric problem (with thin boundary).
3.1. Variant Isoperimetry
@@
where $\Pi(\pi_{X},\pi_{Y})$ is the set of joint probability measures on $\mathcal{X}\times\mathcal{Y}$ with marginals equal to $\pi_{X},\pi_{Y}$ respectively. The existence of the minimizers are well-known; see, e.g., [46, Theorem 1.3]. The minimization problem in (3.4) is called the Monge–Kantorovich’s optimal transport problem [46]. If $\mathcal{X}=\mathcal{Y}$, and $d$ is a complete metric on $\mathcal{X}$ that induces the topology on this space (i.e., $(\mathcal{X},d)$ is a Polish metric space), and $c=d^{p}$ with $p\geq 1$, then $\mathsf{W}_{p}(\pi_{X},\pi_{Y}):=\mathsf{C}(\pi_{X},\pi_{Y})^{1/p}$ is the $p$-th Wasserstein metric between $\pi_{X}$ and $\pi_{Y}$.
@@
Assumption 1: We assume that there is a function $\delta(\epsilon):(0,\infty)\to(0,\infty)$ vanishing as $\epsilon\downarrow 0$ such that
@@
$d^{p}(x^{\prime},y)$ $\geq(d(x,y)-d(x^{\prime},x))^{p}$
$\geq(d(x,y)-\epsilon)^{p}$
(3.9) $\geq\frac{1}{(1+\epsilon)^{p-1}}d^{p}(x,y)-\frac{\epsilon}{1+\epsilon},$
@@
We set $\delta(\epsilon)=\max\{1-\frac{1}{(1+\epsilon)^{p-1}},\frac{\epsilon}{1+\epsilon}\}$. Then, the inequalities in (3.9) imply (3.8).
@@
where $W$ is defined on $\{0,1\}$, and $P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ denotes the set of probability measures concentrated on compact sets. By the convexity of optimal transport cost and the relative entropy, one can easily see that $\tau\mapsto\psi(\alpha,\tau)$ is convex. Define
@@
Without changing the optimal value, the range of $W$ in (3.11) can be changed to any larger but finite set. This point can be proven by following similar proof steps as [48, Theorem 20].
@@
Theorem 8 (Asymptotic Exponent [48, Theorem 12]).
@@
Statement 1 in this theorem follows from the original proof of Statement 1 in [48, Theorem 12] with two slight modifications. Firstly, we replace the power 2 in the equation chain in (56) of [48] by $q=1+\epsilon$, and then apply the Marcinkiewicz–Zygmund inequality to upper bound $\mathbb{E}_{\pi}\left[|c_{n}(x^{n},Y^{n})-n\mu|^{q}\right]$ in terms of $\mathbb{E}_{\pi}\left[c(X,Y)^{q}\right]$ (or $\kappa$). The Marcinkiewicz–Zygmund inequality here refers to that for $1<q<2$, $\mathbb{E}\left[|\sum_{i=1}^{n}X_{i}|^{q}\right]\leq B_{q}\mathbb{E}\left[(\sum_{i=1}^{n}|X_{i}|^{2})^{q/2}\right]\leq B_{q}\sum_{i=1}^{n}\mathbb{E}\left[|X_{i}|^{q}\right]$ for independent $X_{i}$’s such that $\mathbb{E}\left[X_{i}\right]=0$, where $B_{q}$ is a constant only depending on $q$. Secondly, Assumption 1 is modified in the present form in order to include Example 1 (with unbounded metrics). Statement 2 is also given in [48, Theorem 12].
@@
Theorem 8 still holds if the constraints $\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ replaced by $\pi_{XW}\in P(\mathcal{X}\times\{0,1\})$.
@@
$=\{\mathbf{y}:\min_{\sigma}\mathbb{E}_{\mathsf{L}_{\sigma(\mathbf{x}),\mathbf{y}}}[c(X,Y)]\leq\tau,\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$
@@
$\approx\{\mathbf{y}:\mathsf{C}(\pi_{X},\mathsf{L}_{\mathbf{y}})\leq\tau\}$
$=\mathsf{L}^{-1}\{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau\},$
where $\sigma$ denotes a permutation operator and (3.15) is due to Birkhoff’s theorem [46, Page 5]. By Sanov’s theorem again,
@@
By product construction of such $A$’s, we obtain the exponent $\psi(\alpha,\tau)$. That is, empirically $\epsilon$-typical sets are optimal in attaining the optimal exponent.
@@
  • the inf and sup are swapped in (3.19) which follows by the general minimax theorem [37, Theorem 5.2.2] together with the convexity of the relative entropy and the optimal transport cost functional;
@@
  • the inf and sup are swapped in (3.22) which follows by the general minimax theorem [37, Theorem 5.2.2] by identifying that 1) the optimal value of the sup-inf in (3.22) is finite (since upper bounded by $\lambda(\mathsf{C}(\pi_{X|W=w},\nu_{Y})-\tau)$), and 2) by choosing $f,g$ as the zero function, the objective function turns to be $\pi_{Y|W=w}\mapsto D(\pi_{Y|W=w}\|\nu_{Y})-\lambda\tau$ whose sublevels are compact under the weak topology;
@@
$=-\left(\gamma_{Y}^{(\lambda)}(g^{q+1})-\gamma_{Y}^{(\lambda)}(g)\gamma_{Y}^{(\lambda)}(g^{q})\right)$
$\leq 0,$
where the last line follows by the Harris inequality: for any probability measure $\mu$ on $\mathbb{R}$ and increasing functions $f$ and $g$,
$\int_{\mathbb{R}}f(x)g(x)\,{\text{\rm d}}\mu(x)\geq\int_{\mathbb{R}}f(x)\,{\text{\rm d}}\mu(x)\,\int_{\mathbb{R}}g(x)\,{\text{\rm d}}\mu(x).$
Hence, $\gamma_{Y}^{(\lambda)}(g^{q})$ is decreasing in $\lambda\geq 0$. That is, $\gamma_{Y}^{(\lambda)}(g^{q})\leq\gamma_{Y}^{(0)}(g^{q})=\nu_{Y}(g^{q})\leq\nu_{Y}((g^{+})^{q})=:N.$ By the assumption $\nu_{Y}(d(x_{0},\cdot)^{q})<\infty$, it holds that $N<\infty$ since $g^{+}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$. So, $\gamma_{Y}^{(\lambda)}(g^{q})\leq N$ is bounded. Moreover, due to the fact that $g^{-}\leq g$, it holds that $\gamma_{Y}^{(\lambda)}((g^{-})^{q})\leq N$ is also bounded.
Since $g^{-}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$, we have
$g^{-}(y)^{q}\geq(1-\epsilon)d(x_{0},y)^{q}-C$
for some positive constants $\epsilon,C$. Therefore,
@@
(3.25) ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{E}_{\nu_{X}}[e^{\lambda d^{p}(X,x_{0})}]}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}<\infty,\;\forall\lambda\in(0,\infty),}$
@@
$\mathbb{P}\{\mathbf{X}\in A,\,\sum_{i=1}^{n}Z_{i}<n\epsilon^{\prime}\}$
$\geq\mathbb{P}\{\mathbf{X}\in A\}-\mathbb{P}\{\sum_{i=1}^{n}Z_{i}\geq n\epsilon^{\prime}\}$
$\geq e^{-n\alpha}-e^{-nD(\epsilon^{\prime}\|\epsilon)}.$
@@
$\nu_{X}^{\otimes n^{\prime\prime}}(B^{c})$ $=\mathbb{P}\{\sum_{i=n^{\prime}+1}^{n}d^{p}(X_{i},x_{0})>r^{p}\}$
$\leq\mathbb{E}_{\nu}[e^{\lambda d^{p}(X,x_{0})}]^{n^{\prime\prime}}e^{-\lambda r^{p}}$
$=e^{-\lambda\left(r^{p}-n^{\prime\prime}\ell_{\lambda}\right)},$
@@
(3.32) $\lambda\left(\delta^{\prime}-\epsilon^{\prime}\ell_{\lambda}\right)$ $>\beta,$
(3.33) $\mathbb{E}_{\nu}[d^{p}(Y,x_{0})]$ $<\frac{\delta^{\prime}}{\epsilon^{\prime}}.$

### v3

@@ 1.4. Proof Idea and Paper Organization
Ledoux’s semigroup method, as developed in [37, 39], seems insufficient for establishing Theorem 1. This limitation stems from the method’s core strategy: it reduces the isoperimetric problem to the more general noise stability problem. However, obtaining a sufficiently precise solution to the former demands an exact (or sufficiently precise, not merely asymptotic) solution to the latter—an outcome that remains elusive in general. While asymptotically optimal solutions for the noise stability problem are well-established [53, 34, 69, 67], a complete exact resolution has yet to be achieved. Accordingly, a novel approach is required to advance the proof of Theorem 1.
The heuristic reasoning behind our idea is as follows. Given an arbitrary set in a high-dimensional space, its (weighted) surface area does not admit any explicit formula and thus, is difficult to analyze. How can we convert the surface area to a simple formula? A possible way is to tailor the set to a new one that is “regular” in a certain sense, so that the surface area of the resulting set admits a simple formula. However, this is difficult to realize in a direct way, since the surface area is very sensitive to the shape of the set, and thus any subtle change of the set could lead to drastic change in surface area.
Our proof idea of Theorem 1 is as follows, which is new and totally different from the semigroup approach. Firstly, using a generalized Heintze-Karcher theorem [49] from geometric measure theory, we convert the standard isoperimetric problem into a variant version in which the “perimeter” is measured by the size of a thick boundary of the set. Secondly, using the inherently typical subset lemma [2] and its consequence [68] from information theory, we tailor the set into a “regular” one whose thick boundary admits a simple formula. The formula we obtain is expressed in terms of nonlinear transport-entropy inequality. The feasibility of this step is due to the fact that the size of thick boundary is robust to any slight change of the set. Lastly, using nonlinear transport inequality [66, 65] from optimal transport, we simplify the resulting formula and express it in terms of nonlinear log-Sobolev inequality, leading to the equivalence between the standard isoperimetric problem and the nonlinear log-Sobolev inequality.
Our specific strategy is illustrated in Fig. 1.1, which is described in the order opposite to the one given above. Specifically, using tools from optimal transport, in Section 2 we first deduce a specific form of nonlinear transport inequality from the nonlinear log-Sobolev inequality. By tools from information theory, we show in Section 3 that the nonlinear transport inequality in turn implies a variant isoperimetric inequality—the isoperimetric inequality with thick boundaries. Lastly, by tools from geometric measure theory, we show in Section 4 that the variant isoperimetric inequality implies the original isoperimetric inequality. Putting everything together, in Section 5 we deduce that the nonlinear log-Sobolev inequality implies the isoperimetric inequality. As for the other direction, we provide two proofs in Section 6: one is based on information theory and optimal transport, and the other is based on geometric measure theory.
We prove Theorem 2, Theorem 4, Theorem 5, Proposition 2, and Proposition 3 respectively in Appendices A–E.
@@ 2. From Log-Sobolev to Transport
2.1. From Log-Sobolev to Transport
@@
It is well known that Sobolev-type inequalities (including log-Sobolev inequalities) imply Talagrand’s transport inequalities; this point was shown by Otto and Villani [54]. A nonlinear analogue of this result was given by F. Y. Wang [66, Theorem 2.1] which is proven by extending Otto and Villani’s idea. Specifically, F. Y. Wang showed that if
(2.1) $\theta(D(\lambda\|\nu))\leq I(\lambda\|\nu),\;\forall\lambda\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
for some function $\theta$, then
(2.2) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\int_{0}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r,\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}).$
In the present paper, we require a variant of F. Y. Wang’s result.
Theorem 6 (From Nonlinear Log-Sobolev to Nonlinear Transport ).
Assume that (2.1) holds for some function $\theta$. Then, for any $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ and any $0\leq s\leq D(\mu\|\nu)$, there is a probability measure $\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\pi\|\nu)=s$ and
Proof.
Define a Laplace-type operator $L=\Delta-\nabla V\cdot\nabla$. Let $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ be a probability measure with density $h=\frac{{\text{\rm d}}\mu}{{\text{\rm d}}\nu}$ such that
@@
The heat equation $\partial_{t}\rho=L\rho$ with initial data $\rho_{0}=h$ admits a solution $\rho_{t}$; see [54]. Let $\mu_{t}=\rho_{t}\nu$. The heat equation has intimate relations with the entropy and the Wasserstein metric in the sense that the heat equation is the gradient flow of the entropy functional $D$ with respect to the Wasserstein metric $\mathsf{W}_{2}$.
By an approximation argument in [54], for every $\mu\in P_{2}^{\mathrm{ac}}(M)$, there is a sequence of probability measures $\mu_{k}$ such that the density $h_{k}:={\text{\rm d}}\mu_{k}/{\text{\rm d}}\nu$ for each $k$ satisfies (2.4) and moreover, as $k\to\infty$,
@@
Given any $0<s<D(\mu\|\nu)$, it holds that $0<s<D(\mu_{k}\|\nu)$ for all sufficiently large $k$. Following the proof of F. Y. Wang [66, Theorem 2.1] (taking the integration $\int_{t}^{\infty}$ for a proper $t$, instead of $\int_{0}^{\infty}$ [66, (2.5)]), it is obtained that for each such $k$, there is a probability measure $\pi_{k}$ (the heat flow at $t$ with initial data $\mu_{k}$) such that $D(\pi_{k}\|\nu)=s$ and
@@
Note that (2.6) implies that $\mu_{k}$ belongs to a sublevel set of the relative entropy $\pi\mapsto D(\pi\|\nu)$, and so does $\pi_{k}$. Since the sublevel set of the relative entropy is weakly compact, $\pi_{k}$ converges to some $\pi^{*}$ weakly. Since $\mu$ has finite second moment, (2.5) implies that the sequence $\mu_{k}$ has bounded second moment. In this case, $\pi_{k}\to\pi^{*}$ weakly is equivalent to $\mathsf{W}_{2}\left(\pi_{k},\pi^{*}\right)\to 0$. Taking limits for (2.7) and applying the triangle inequality yield
@@
The importance of Theorem 6 is as follows. Given $\tau\geq 0,\mu,\nu$, consider the following optimization problem:
@@ 2.2. Generalization to $\Phi$-Sobolev
Theorem 6 can be generalized to the general $\Phi$-entropy setting (also known as $U$-functional or $f$-divergence). Let $\Phi:[0,+\infty)\rightarrow\mathbb{R}$ be a continuous convex function, twice differentiable on $(0,+\infty)$. The $\Phi$-entropy is defined as
@@
A common requirement for the function $\Phi$ is the following, which guarantees the displacement convexity of the $\Phi$-entropy [65, Chapter 16].
@@ 3. From Transport to Variant Isoperimetry
3.1. From Transport to Variant Isoperimetry
@@
where $\Pi(\pi_{X},\pi_{Y})$ is the set of joint probability measures on $\mathcal{X}\times\mathcal{Y}$ with marginals equal to $\pi_{X},\pi_{Y}$ respectively. The existence of the minimizers are well-known; see, e.g., [64, Theorem 1.3]. The minimization problem in (3.4) is called the Monge–Kantorovich’s optimal transport problem [64]. If $\mathcal{X}=\mathcal{Y}$, and $d$ is a complete metric on $\mathcal{X}$ that induces the topology on this space (i.e., $(\mathcal{X},d)$ is a Polish metric space), and $c=d^{p}$ with $p\geq 1$, then $\mathsf{W}_{p}(\pi_{X},\pi_{Y}):=\mathsf{C}(\pi_{X},\pi_{Y})^{1/p}$ is the $p$-th Wasserstein metric between $\pi_{X}$ and $\pi_{Y}$.
@@
Assumption 3.
We assume that there is a function $\delta(\epsilon):(0,\infty)\to(0,\infty)$ vanishing as $\epsilon\downarrow 0$ such that
@@
(3.9) $d^{p}(x^{\prime},y)$ $\geq(d(x,y)-d(x^{\prime},x))^{p}\geq(d(x,y)-\epsilon)^{p}\geq\frac{1}{(1+\epsilon)^{p-1}}d^{p}(x,y)-\epsilon,$
@@
We set $\delta(\epsilon)=\max\{1-\frac{1}{(1+\epsilon)^{p-1}},\epsilon\}$. Then, the inequalities in (3.9) imply (3.8).
@@
where $W$ is defined on $\{0,1\}$, and $P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ denotes the set of probability measures concentrated on compact sets. The infimum above actually represents a nonlinear transport inequality. By the convexity of optimal transport cost and the relative entropy, one can easily see that $\tau\mapsto\psi(\alpha,\tau)$ is convex. Define
@@
Without changing the optimal value, the range of $W$ in (3.11) can be changed to any larger but finite set. This point can be proven by following similar proof steps as [68, Theorem 20].
@@
Theorem 8 (Asymptotic Exponent [68, Theorem 12]).
@@
Statement 1 in this theorem follows from the original proof of Statement 1 in [68, Theorem 12] with two slight modifications. Firstly, we replace the power 2 in the equation chain in (56) of [68] by $q=1+\epsilon$, and then apply the Marcinkiewicz–Zygmund inequality to upper bound $\mathbb{E}_{\pi}\left[|c_{n}(x^{n},Y^{n})-n\mu|^{q}\right]$ in terms of $\mathbb{E}_{\pi}\left[c(X,Y)^{q}\right]$ (or $\kappa$). The Marcinkiewicz–Zygmund inequality here refers to that for $1<q<2$, $\mathbb{E}\left[|\sum_{i=1}^{n}X_{i}|^{q}\right]\leq B_{q}\mathbb{E}\left[(\sum_{i=1}^{n}|X_{i}|^{2})^{q/2}\right]\leq B_{q}\sum_{i=1}^{n}\mathbb{E}\left[|X_{i}|^{q}\right]$ for independent $X_{i}$’s such that $\mathbb{E}\left[X_{i}\right]=0$, where $B_{q}$ is a constant only depending on $q$. Secondly, Assumption 3 is modified in the present form in order to include Example 1 (with unbounded metrics). Statement 2 is also given in [68, Theorem 12].
@@
By checking the proof, Theorem 8 still holds if the constraint $\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ is replaced by $\pi_{XW}\in P(\mathcal{X}\times\{0,1\})$.
The inherently typical subset lemma in [2] plays a key role in the proof of the theorem above, which enables us to tailor a set into a new one that is “regular” in some sense, so that we can analyze the asymptotic expression of the boundary size of the set.
@@
$\approx\mathsf{L}^{-1}\{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau\},$
where (3.15) is due to Birkhoff’s theorem, $\min_{\sigma}\mathbb{E}_{\mathsf{L}_{\sigma(\mathbf{x}),\mathbf{y}}}[c(X,Y)]=\mathsf{C}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})$, with $\sigma$ denoting a permutation operator [64, Page 5]. By Sanov’s theorem again,
@@
By product construction of such $A$’s (forming a conditional empirically typical set), we obtain the exponent $\psi(\alpha,\tau)$. That is, conditional empirically $\epsilon$-typical sets are optimal in attaining the optimal exponent.
@@
  • the inf and sup are swapped in (3.19) which follows by the general minimax theorem [52, Theorem 5.2.2] together with the convexity of the relative entropy and the optimal transport cost functional;
@@
  • the inf and sup are swapped in (3.22) which follows by the general minimax theorem [52, Theorem 5.2.2] by identifying that 1) the optimal value of the sup-inf in (3.22) is finite (since upper bounded by $\lambda(\mathsf{C}(\pi_{X|W=w},\nu_{Y})-\tau)$), and 2) by choosing $f,g$ as the zero function, the objective function turns to be $\pi_{Y|W=w}\mapsto D(\pi_{Y|W=w}\|\nu_{Y})-\lambda\tau$ whose sublevels are compact under the weak topology;
@@
$=-\left(\gamma_{Y}^{(\lambda)}(g^{q+1})-\gamma_{Y}^{(\lambda)}(g)\gamma_{Y}^{(\lambda)}(g^{q})\right)\leq 0,$
where the last inequality follows by the Harris inequality for the random variable $g(Y)$ with $Y\sim\gamma_{Y}^{(\lambda)}$ and increasing functions $t\mapsto t$ and $t\mapsto t^{q}$. Hence, $\gamma_{Y}^{(\lambda)}(g^{q})$ is decreasing in $\lambda\geq 0$. That is, $\gamma_{Y}^{(\lambda)}(g^{q})\leq\gamma_{Y}^{(0)}(g^{q})=\nu_{Y}(g^{q})\leq\nu_{Y}((g^{+})^{q})=:N.$ By the assumption $\nu_{Y}(d(x_{0},\cdot)^{q})<\infty$, it holds that $N<\infty$ since $g^{+}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$. So, $\gamma_{Y}^{(\lambda)}(g^{q})\leq N$ is bounded. Moreover, due to the fact that $g^{-}\leq g$, it holds that $\gamma_{Y}^{(\lambda)}((g^{-})^{q})\leq N$ is also bounded.
Since $g^{-}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$, we have $g^{-}(y)^{q}\geq(1-\epsilon)d(x_{0},y)^{q}-C$ for some positive constants $\epsilon,C$. Therefore,
@@
(3.25) $\mathbb{E}_{\nu_{X}}[e^{\lambda d^{p}(X,x_{0})}]$ $<\infty,\;\forall\lambda\in(0,\infty),$
@@
$\mathbb{P}\{\mathbf{X}\in A,\,\sum_{i=1}^{n}Z_{i}<n\epsilon^{\prime}\}$ $\geq\mathbb{P}\{\mathbf{X}\in A\}-\mathbb{P}\{\sum_{i=1}^{n}Z_{i}\geq n\epsilon^{\prime}\}\geq e^{-n\alpha}-e^{-nD(\epsilon^{\prime}\|\epsilon)}.$
@@
$\nu_{X}^{\otimes n^{\prime\prime}}(B^{c})$ $=\mathbb{P}\{\sum_{i=n^{\prime}+1}^{n}d^{p}(X_{i},x_{0})>r^{p}\}\leq\mathbb{E}_{\nu}[e^{\lambda d^{p}(X,x_{0})}]^{n^{\prime\prime}}e^{-\lambda r^{p}}=e^{-\lambda\left(r^{p}-n^{\prime\prime}\ell_{\lambda}\right)},$
@@
(3.32) $\lambda\left(\delta^{\prime}-\epsilon^{\prime}\ell_{\lambda}\right)$ $>\beta\textrm{ and }\mathbb{E}_{\nu}[d^{p}(Y,x_{0})]<\frac{\delta^{\prime}}{\epsilon^{\prime}}.$